### v2

# Statistical Complexity of Heterogeneous Geometric Networks

Abstract
Heterogeneity and geometry are key explanatory components underlying the structure of real-world networks. The relationship between these components and the statistical complexity of networks is not well understood. We introduce a parsimonious normalised measure of statistical complexity for networks. The measure is trivially 0 in regular graphs and we prove that this measure tends to 0 in Erdös-Rényi random graphs in the thermodynamic limit. We go on to demonstrate that greater complexity arises from the combination of hierarchical and geometric components to the network structure than either on their own. Further, the levels of complexity achieved are similar to those found in many real-world networks. We also find that real world networks establish connections in a way which increases complexity and which our null models and a range of attachment mechanisms fail to explain. This underlines the non-trivial nature of statistical complexity in real-world networks and provides foundations for the comparative analysis of network complexity within and across disciplines.
I Introduction
[…]
Yet, while randomly generated information may be difficult to deterministically reproduce, it is not structurally complex in a statistical sense. Indeed, the statistical properties of randomly generated information are defined a-priori and are evidently simple. This led the field of dynamical systems to lay out a different conceptual framework of complexity. In this view, a measure of complexity should go to zero for regular and random structures in the thermodynamic limit (as number of components goes to infinity), while being higher for systems presenting non-trivial and diverse correlations [4, 5]. One particularly important point of developing measures of statistical complexity is that using a scale between regular and random with complexity somewhere in the middle, a common approach from an information theoretic angle, does not allow for a useful measure of complexity itself [4, 5]. Instead, we need a scale between the simple (regularity and randomness both having uniform generational principles) and the complex, allowing us to directly measure the extent of complexity in any given system.
When it comes to studying complexity in networks, we are concerned with the complexity of interactions– essentially, how diverse the connectivity patterns in the network are. While others borrow from the algorithmic view of complexity [6, 7], here we are concerned with the statistical complexity of networks. A notable early work on statistical complexity of networks introduced a measure called the network diversity score and provided a comprehensive overview of the problem [8]. Another work considered statistical complexity in networks from an information theoretic angle, multiplying Jensen Shannon divergence of a network with network entropy [9]. In this study, we establish a normalisation of Hierarchical Complexity (HC) as a network statistical complexity measure. In contrast to the network diversity score, HC is a parsimonious calculation of a single feature of a network which can be applied to individual networks and does not require a-priori categorisation of the complexity of the networks in question [10]. In contrast to the measure in [9] it does not require a reference graph, displays strong independence to network density and vanishes in the thermodynamic limit for Erdös-Rényi random graphs.
The hierarchy referred to here is the degree hierarchy of the network. A hierarchically complex system is one for which diversity of connectivity patterns are found across hierarchical levels (either individual degrees or ranges of degrees called tiers). Its introduction was motivated by the need to measure the complicated hierarchical networks of brain function and structure where it was expected that diverse functionality would be reflected in diverse connectivity patterns. HC has so far seen limited application in fairly small ($n<100$) macro-scale human brain networks [10, 11, 12, 13, 14] and a corpus of real-world networks of varied origin ($n<5000$) [15].
A major issue with the generalisability of the HC measure is that it is not normalised by number of nodes or number of edges. Two similarly derived networks with different numbers of nodes or edges can be expected to have different values of HC. This paper addresses this issue by introducing a normalised HC measure. We show mathematically that this measure is bounded above by $2$ and satisfies the statistical complexity definition of being asymptotically zero for Erdös-Rényi random graphs (an appropriate equivalent to randomness in dynamical systems). We then go on to explore results of this normalisation on different types of random graphs with the two most evident structural properties relevant to real-world networks, hierarchy and geometry. Hierarchy here relates to the distribution of node fitness [16] or popularity [17]. Geometry relates to the latent space of similarities between nodes [18, 19]. Combining hierarchy and geometry successfully captures many of the properties of real-world networks [20], but whether these properties are enough to explain the statistical complexity of networks is not known. After this we go on to explore normalised HC in real-world networks. Here, an unexpected relationship between HC and density is noted. Finally, we explore explanations for this relationship by applying different kinds of attachment mechanisms based on degrees and overlap of node neighbourhoods. Combined, the results reveal non-trivial hierarchical complexity in real-world networks and we open the way for more reliable and robust applications of hierarchical complexity across network domains.
II Theory
Key themes within this work are encapsulated within the image in Fig. 1 which illustrates hierarchical complexity arising from a combination of geometric and hierarchical structure. Henceforth, unless specified otherwise, let $G$ be a graph with $n$ nodes, $m$ edges and density $d$.
Figure 1: On the left we see a geometric graph with a regular structure. Node shapes indicate distinct degrees while colours indicate distinct, repeating neighbourhood degree sequences. On the right, nodes are randomly assigned different numbers of connections. These connections are made to the closest nodes, maintaining a geometric nature, but now we see the diversity of structure this opens up. Again, different shapes indicate distinct degrees, but now there are many unique neighbourhood degree sequences which remain colourless. This diversity reflects a higher hierarchical complexity.
II-A Hierarchical complexity
To compute hierarchical complexity, we first define the Neighbourhood Degree Sequence (NDS) of a node $i$ of degree $k$ as
$s_{i}=\{s_{i1},s_{i2},\dots,s_{ik}\}$ (1)
where the $s_{ij}$’s are the degree of the nodes to which $i$ is connected such that $s_{i1}\leq s_{i2},\dots,\leq s_{ik}$. Then, for all $\ell$ nodes of degree $k$, we can stack their NDSs into an ×$\ell\times k$ matrix:
$\mathbf{S}_{k}(G)=\begin{bmatrix}s_{11}&s_{12}&\dots&s_{1k}\\ s_{21}&s_{22}&\dots&s_{2k}\\ \vdots&\vdots&\ddots&\vdots\\ s_{l1}&s_{l2}&\dots&s_{lk}\end{bmatrix}$ (2)
The original definition of hierarchical complexity for degree $k$ takes the variance over the columns of this matrix and then averages across columns:
$R_{k}(G)=\frac{\sum_{j=1}^{k}\sigma_{j}^{2}}{k}$ (3)
where $\sigma_{j}^{2}$ is the variance of the $j$th column of $\mathbf{S}_{k}(G)$. The global measure is then the average of this over degrees:
$R(G)=\frac{1}{|\mathcal{D}_{2}|}\sum_{k\in\mathcal{D}_{2}}R_{k}(G).$ (4)
where $\mathcal{D}_{2}$ is the set of degrees in the graph taken by at least two nodes (since variance is only meaningful over at least two elements).
II-B Normalised Hierarchical Complexity
It is clear that the above definition of hierarchical complexity depends on network size. Networks of larger size have greater potential for larger degrees, which will influence the variances within equation (3). Indeed, the maximum variance of numbers in $[1,n-1]$ can occur with the sample $\{1,n-1\}$ (or any equal number of 1s and $n-1$s) which has variance $((n-2)/2)^{2}$. Furthermore, computations have demonstrated that hierarchical complexity also correlates with number of edges.
Lack of normalisation to number of nodes and/or density is not unusual in network science. Indeed, the latter is essentially common place– consider two of the most widely considered network metrics, the global clustering coefficient and global efficiency which are both maximised in complete graphs.
It is then of note that we can propose the following as a normalised measure of hierarchical complexity to both $n$ and $d$.
Definition 1.
Let $G$ be a network, we define the $k$-th normalised hierarchical complexity as:
$\hat{R}_{k}(G)=\begin{dcases}\frac{\sum_{j=1}^{k}\sigma_{j}}{(1-d)m},&\mbox{ if $m\not=0$ and $d\not=1$}\\ 0,&\mbox{ if $m=0$ or $d=1$}\\ \end{dcases}$ (5)
where $d$ is the density of $G$, $m$ is the number of edges in $G$, and $\sigma_{j}$ is the standard deviation of the $j$th column of $\mathbf{S}_{k}(G)$, that is, the matrix where each row is the ordered degree sequence of a node of degree k. Thus we propose the normalised global measure as
$\hat{R}(G)=\frac{1}{|\mathcal{D}_{2}|}\sum_{k\in\mathcal{D}_{2}}\hat{R}_{k}(G).$ (6)
To justify this normalisation, we note that the normalised hierarchical complexity is bounded.
Theorem 1.
The normalised hierarchical complexity is bounded above by $2$, that is, $\hat{R}(G)\leq 2$ for every graph $G$.
Proof.
Consider $\hat{R}_{k}$. Let $d_{j}$ be the difference of the max and min elements of the $j$-th column of $\mathbf{S}_{k}(G)$, so $\sigma_{j}\leq\frac{d_{j}}{2}$. Let $\alpha=\sum_{j=1}^{k}d_{j}$.
Now, let $x_{j}$ be the maximum degree of the $j$’th column of $\mathbf{S}_{k}(G)$. We claim that $2m\geq\sum_{j=1}^{k}x_{j}$. If $x_{1},\ldots,x_{k}$ are degrees of distinct vertices, this follows immediately, but it could occur that $x_{i}$ and $x_{j}$ correspond to the same vertex. Note that $x_{1}\geq x_{2}\geq\dots\geq x_{k}$, since the rows are ordered in decreasing order. Also note that if $x_{i}=x_{j}$, with $i<j$, and $x_{j}$ is in row $r$, then the $i$-th entry in row $r$ must also equal $x_{i}$, since it cannot be smaller than something to it’s right $x_{j}$, and cannot be bigger than the max of it’s column $x_{i}$. And since the degrees in each row correspond to distinct vertices (as they are given by the neighourhood), if $x_{i}=x_{j}$ then there are at least two vertices of degree $x_{i}$. And generalising this, if $x_{i}$ appears as the degree of the same vertex $\ell$ times, then we can find $\ell$ distinct vertices of the same degree, so $2m\geq\sum_{j=1}^{k}x_{j}$. Also note that $x_{j}\geq d_{j}$, for all $j$, therefore
$2m\geq\sum_{j=1}^{k}x_{j}\geq\sum_{j=1}^{k}d_{j}=\alpha.$
Considering the minimal elements of each column and applying a similar argument to above we can also deduce that $2m^{\prime}\geq\alpha$, where $m^{\prime}=\frac{n(n-1)}{2}-m$ is the number of non-edges. Therefore, $\min(m,m^{\prime})\geq\alpha$. Moreover, $\max(m,m^{\prime})\geq\frac{n(n-1)}{4}$, since either at least half the edges are there or half are not.
So we can bound $(1-d)m$ below by:
$(1-d)m$ $=\frac{2mm^{\prime}}{n(n-1)}=\frac{2\max(m,m^{\prime})\min(m,m^{\prime})}{n(n-1)}$
$\geq\frac{(\frac{n(n-1)}{4})\alpha)}{n(n-1)}=\frac{\alpha}{4}$
Therefore,
$\hat{R}_{k}(G)=\frac{\sum_{j=1}^{k}\sigma_{j}}{(1-d)m}\leq\frac{\frac{\alpha}{2}}{(1-d)m}\leq\frac{\frac{\alpha}{2}}{\frac{\alpha}{4}}=2$
So $\hat{R}_{k}(G)$ is bounded above by $2$ for all $k$, and $\hat{R}(G)$ is the average of values bounded above by $2$, hence $\hat{R}(G)\leq 2$. ∎
There are several things to note about the formula (5). Firstly, instead of variance across neighbourhood degree sequences, we here opt for standard deviation. The distribution of the standard deviation over multiple samples will in general be more symmetric than that of the variance which will be right-skewed and standard deviation is generally a more appropriate measure when normalising.
Secondly, the division by $(1-d)m$ is the term which acts to normalise the measure. This term was borrowed from the normalisation of degree variance [21]. There, it was shown to normalise degree variance and bound it below 1 for all graphs.
Thirdly, instead of taking the mean over the standard deviations this normalisation takes the sum. In actuality we can consider this as a multiplication of the mean by the degree of the neighbourhood degree sequences $k$ (cancelling out the $k$ on the denominator of the average). This effectively takes account of the sampling error of taking the mean over the $\sigma_{j}$’s. The sampling error of the mean over $k$ samples is:
$\frac{\sigma}{\sqrt{k}},$ (7)
where $\sigma$ here is the standard deviation of the $k$ element-wise standard deviations, so that the accuracy of the mean depends on the degree $k$.
Note that in the normalisation presented we do not multiply by $\sqrt{k}$ to standardise these measurements, but by $k$ itself. This is because it is also linked with the division by $m$.
We do not believe the bound of $2$ given in Theorem 1 is tight. Through considered construction of a disconnected graph which exploits variances of 1 degree nodes, we find the largest value for $\hat{R}$ tends to $\frac{1}{3}$, which occurs with the following graph family:
Example 1.
Consider the graph:
$\dots$$\dots$$\dots$
where each column of nodes consists of $\frac{n-1}{4}$ nodes.
Note that $R_{1}=\frac{n-3}{4}$, no other degree contributes to $R$, $m=\frac{3(n-1)}{4}$ and $d=\frac{3}{2n}$. So
$\hat{R}=\hat{R}_{1}=\frac{\frac{n-3}{4}}{(1-\frac{3}{2n})\frac{3(n-1)}{4}}\rightarrow\frac{1}{3},\,\,\,\,\,\,\,\text{ as }n\rightarrow\infty$
We believe that the above family of graphs gives the largest value for $\hat{R}$, but we leave this as a conjecture:
Conjecture 1.
For any graph $G$ we have $\hat{R}(G)<\frac{1}{3}$.
While we in no way claim the above family of graphs is statistically or otherwise complex, the intended application of the measure is for connected graphs with many different degrees making a contribution. We can ensure this edge case goes to zero by using the corrective term for multiply-ordered degrees as described in [15], but for most intended purposes this is not necessary.
II-C Expected Complexity Values
Next we show that $\hat{R}$ satisfies the conditions of a statistical complexity measure of being 0 in the thermodynamic limit for the Erdös-Rényi random graph. It is known (see [10]) that if a graph is regular, that is, every node has the same degree, then the hierarchical complexity is 00– since we have $n$ nodes of degree $k$ all with neighbourhood degree sequences $\{k,k,...,k\}$.
We can also show that for the other end of the entropic spectrum, Erdös-Rényi random graphs, that the hierarchical complexity tends to 0, as $n$ tends to infinity. To do so we need a formula for the standard deviation of the $i$-th largest sample from this distribution, this is known as the $i$-th order statistic, see [22] for background on order statistics. We can use known results on order statistics to derive the following:
Theorem 2.
Fix $d\in[0,1]$. The normalised hierarchical complexity of an Erdös-Rényi graph $\operatorname*{ER}(n,d)$ tends to 00 as $n$ tends to infinity, that is,
$\lim_{n\rightarrow\infty}\hat{R}(\operatorname*{ER}(n,d))=0.$
Proof.
For brevity let $\hat{R}_{n,d}:=\hat{R}(\operatorname*{ER}(n,d))$. Note that if $d=0$ or $d=1$, then $\operatorname*{ER}(n,d)$ is regular, thus $\hat{R}_{n,d}=0$ for all $n$, and the result holds. So assume $d\in(0,1)$. We begin by giving an approximation of $\hat{R}_{n,d}$.
First note that the node degrees of an Erdös-Rényi graph are sampled from the binomial distribution ${B(n-1,d)}$. Whilst this sampling is not strictly independent, the dependence is very weak, and the correlation tends to zero as $n$ tends to infinity [23], and thus can be disregarded in our approximation and asymptotic analysis.
In [24, Theorem 9.1] an approximate formula for the standard deviation of the $i$-th order statistic of $k$ samples for a continuous distribution with PDF $\phi(x)$ and CDF $\Phi(X)$ is given by:
$\sigma_{i}\approx\frac{1}{\phi(\Phi^{-1}(\frac{i}{k+1}))}\sqrt{\frac{i(k-i+1)}{(k+1)^{2}(k+2)}}.$ (8)
By [25, Equation (1.1)] this approximation also holds for discrete distributions.
So to get the $k$’th normalised hierarchical complexity of $\operatorname*{ER}(n,d)$, we let $X$ be the binomial distribution ${B(n-1,d)}$, sum (8) across $i=1,\ldots,k$ and divide by our normalisation, which gives:
$\hat{R}_{k}(\operatorname*{ER}(n,d))$ $\approx\frac{\sum_{i=1}^{k}\frac{1}{\phi(\Phi^{-1}(\frac{i}{k+1}))}\sqrt{\frac{i(k-i+1)}{(k+1)^{2}(k+2)}}}{(1-d)m}$
$=\frac{2\sum_{i=1}^{k}\frac{\sqrt{i(k-i+1)}}{\phi(\Phi^{-1}(\frac{i}{k+1}))}}{d(1-d)n(n-1)(k+1)\sqrt{k+2}}.$
The second equality is because the expected number of edges is $m=\frac{n(n-1)d}{2}$.
We can approximate the expected smallest and largest degree in $\operatorname*{ER}(n,d)$, using [Equation 4.5.1][22] which gives an approximation for the $i$’th order statistic as $\Phi^{-1}(\frac{i}{n+1})$, for sufficiently large $n$. This gives our lower and upper summands $a$ and $b$, and the global formula is given by averaging $\hat{R}_{k}$ between $a$ and $b$.
So for large $n$ the global hierarchical complexity can be approximated by
$\hat{R}(\operatorname*{ER}(n,d))\approx\frac{2\sum_{k=a}^{b}\frac{\sum_{i=1}^{k}\frac{\sqrt{i(k-i+1)}}{\phi(\Phi^{-1}(\frac{i}{k+1}))}}{(k+1)\sqrt{k+2}}}{(b-a)d(1-d)(n-1)n},$ (9)
where $\phi$ and $\Phi$ are the PMF and CDF, respectively, of the binomial distribution $B(n-1,d)$, and $a=\lfloor\Phi^{-1}(\frac{1}{n})\rfloor$ and $b=\lceil\Phi^{-1}(\frac{n-1}{n})\rceil$. Note that this approximation is not particularly close, particularly for small $n$, but it is sufficient to consider the limit of the complexity as $n$ grows.
The binomial distribution can be approximated by the normal distribution, for which the quantile function $\Phi^{-1}$ is known in terms of the inverse error function. Combining this with an approximation of the inverse error function [26, Equation 13] we get:
$\Phi^{-1}(x)$ $\approx((n-1)d+\sqrt{2(n-1)d(1-d)}\text{ erf}^{-1}(2x-1)$
$\approx(n-1)d+\sqrt{-\ln(4x(1-x))2(n-1)d(1-d)}$ (10)
Combining Equation (10) with the De Moivre-Laplace approximation of the binomial PMF we get:
$\phi$ $(\Phi^{-1}(x))\approx\frac{\exp\left(-\frac{(\Phi^{-1}(x)-(n-1)d)^{2}}{2(n-1)d(1-d)}\right)}{\sqrt{2\pi(n-1)d(1-d)}}$
$\approx\frac{\exp\left(-\frac{((n-1)d+\sqrt{-\ln(4x(1-x))2(n-1)d(1-d)}-(n-1)d)^{2}}{2(n-1)d(1-d)}\right)}{\sqrt{2\pi(n-1)d(1-d)}}$
$=\frac{\exp\left(\ln(4x(1-x))\right)}{\sqrt{2\pi(n-1)d(1-d)}}=\frac{4x(1-x)}{\sqrt{2\pi(n-1)d(1-d)}}$ (11)
We can then use Equation (11) to show that asymptotically $\hat{R}_{n,d}$ is bounded above by zero:
$\hat{R}_{n,d}$ $\approx\frac{2}{(b-a)d(1-d)n(n-1)}\sum_{k=a}^{b}\frac{\sum_{i=1}^{k}\frac{\sqrt{i(k-i+1)}}{\phi(\Phi^{-1}(\frac{i}{k+1}))}}{(k+1)\sqrt{k+2}}$
$\approx\sum_{k=a}^{b}\frac{2\sum_{i=1}^{k}\frac{\sqrt{2\pi(n-1)d(1-d)}\sqrt{i(k-i+1)}}{4\frac{i}{k+1}(1-\frac{i}{k+1})}}{(b-a)d(1-d)n(n-1)(k+1)\sqrt{k+2}}$
$=\sum_{k=a}^{b}\frac{\sqrt{\pi}\sum_{i=1}^{k}\frac{k+1}{\sqrt{i(k-i+1)}}}{n(b-a)\sqrt{2d(1-d)(n-1)(k+2)}}$
$=\frac{\sqrt{\pi}\sum_{k=a}^{b}\sum_{i=1}^{k}\frac{k+1}{\sqrt{i(k-i+1)(k+2)}}}{n(b-a)\sqrt{2d(1-d)(n-1)}}$
$\leq\frac{\sqrt{\pi}\sum_{k=a}^{b}\sum_{i=1}^{k}1}{n(b-a)\sqrt{2d(1-d)(n-1)}}$
$\leq\frac{\sqrt{\pi}(b-a)b}{n(b-a)\sqrt{2d(1-d)(n-1)}}\xrightarrow[n\rightarrow\infty]{}0$ (12)
The final step follows since $d$ is fixed and $b$ is always smaller than $n$, thus the denominator grows at least $\sqrt{n-1}$ faster than the numerator.
We conjecture that a similar result holds for random geometric graph (RGG) (see Section III-A for the definition of RGG’s) with a fixed average degree of $b$ (note in this case, $d\rightarrow 0$ as $n\rightarrow\infty$). In particular, when we randomly position $n$ nodes on the unitary Euclidean plane and connect two nodes whenever they are within radius $r=\sqrt{b/(n\pi)}$ of each other. This radius $r$ is selected to ensure we obtain a graph with the required average degree. In this version of an RGG, the degree distribution is also the binomial distribution (see [27]), so we can apply a similar technique as used in the proof of Theorem 2. However, RGGs have non-trivial degree correlations violating statistical assumptions used in the Erdös-Rényi case [27]. Nevertheless, we conjecture that a different bound can be obtained that does tend to 00, we leave this as an open problem.
Conjecture 2.
Let $RGG(n,r)$ be the random geometric graph in the unitary plane with $n$ nodes and radius $r=\sqrt{\frac{b}{n\pi}}$, for a fixed $b\in\mathbb{N}$. We have
$\lim_{n\rightarrow\infty}\hat{R}(RGG(n,r))=0.$
We’ve seen that the formula for hierarchical complexity is closely linked to order statistics and the quantile function of probability distributions. Very few closed formulas exist for quantile functions, and they are known to be difficult to analyse. Due to this we are otherwise limited in our analytical treatment of this measure. The remainder of the paper shows the validity of this normalisation in application and we use it to derive novel insights from models and real data. Particularly, we pursue the hypothesis that statistical complexity arises naturally through a combination of hierarchical and geometric components to network connectivity.
III Methods
III-A Network models
Our understanding of this normalisation is aided by its application to different network models and studying the behaviour of our normalisation of HC as we change the size and density of the network. In the following, a graph refers to a mathematical object of a set of nodes with adjoining edges. A network refers to a graph representation of the relationships or connections between components of a real-world complex system.
Firstly, we use Erdös-Rényi (E-R) Random Graphs. E-R random graphs are generated using random uniform edge probabilities in $[0,1]$ [28]. They give an indication of the behaviour of an ‘average’ graph of a given size and density. That being said, they do not give any indication of the behaviour of an ‘average’ network as it lacks many of the basic characteristics common to networks such as a relatively high clustering coefficient and degree heteregoneity.
Random Geometric Graphs (RGG) are generated from randomly sampled co-ordinates in the unit cube (i.e. 3D) [29]. These samples then represent nodes and the inverse distances between node pairs are the weights of the edges between them. For a desired network density, we select the $m$ largest weights as our graph edges. RGGs have properties of high clustering desirable for networks, however they also lack the characteristic degree heterogeneity of networks.
Surface-Depth (S-D) models provide geometric graphs with heterogeneous degree distributions which show distinct similarities to many real-world networks [20]. We shall refer to these models throughout as Random Hierarchical Geometric Graphs (RHGG) since they combine a geometric component to connections with a hierarchical component, both of which we also want to study in isolation. These are generated using two parameters, the $\sigma$ of a log-normal distribution and the number of dimensions, $q$, of a random geometric graph. The weights of the edges are then defined as
$w_{ij}=d_{ij}(s_{i}+s_{j}),$ (13)
where $d_{ij}$ is the inverse distance between nodes $i$ and $j$ in a random geometric graph with $q$ dimensions and each $s_{i}$ is a random sample from a log-normal distribution $LN(\mu,\sigma)$. Again, for a desired network density, we select the $m$ largest weights as our graph edges. In this study, we perform a basic exploration with $q=3$, $\mu=0$, and $\sigma$ fixed at 0.2 as this produces graphs with a suitable heterogeneity. We then also explore the effect of varying $\sigma$, and so heterogeneity of geometric graphs, on hierarchical complexity.
Configuration models of these RHGGs allow us to probe the extent to which hierarchical complexity of the RHGGs can be attributed merely to the hierarchical structure of the network. We refer to these as Random Heterogeneous Graphs (RHGs) to emphasise the relationship with our other models. Briefly, configuration models work by fixing the degree distribution of a network but otherwise randomising the connections. Each node is provided with a number of stubs equal to its degree. These stubs are then randomly paired between nodes to establish edges [30].
III-B Attachment mechanisms
We used four different attachment mechanisms to observe the effect on hierarchical complexity of increasing density in different ways on real-world networks.
For all node pairs without edges, $(i,j)\notin\mathcal{E}$ we considered:
Random attachment: edge probabilities are uniform.
Hierarchical attachment: similar in fashion to preferential attachment for addition of new nodes to a network, the probability of a new edge is proportional to the sum of the degrees of the nodes:
$p_{ij}\sim k_{i}+k_{j}.$ (14)
Similarity attachment: for $g_{i},g_{j}$ the neighbourhoods of nodes $i$ and $j$, the probability of edge $(i,j)$ occurring is proportional to the Jaccard index of their neighbourhoods:
$p_{ij}\sim J(g_{i},g_{j})=\frac{|g_{i}\cap g_{j}|}{|g_{i}\cup g_{j}|}.$ (15)
Combined hierarchical and similarity attachment: the probability of a new edge is proportional to the intersection of their neighbourhoods, essentially removing the size normalisation of the Jaccard index:
$p_{ij}\sim|g_{i}\cap g_{j}|$ (16)
so that the probability of connection is dependent on the size of the neighbourhoods and the overlap of the neighbourhoods.
To get the probabilities, $p_{ij}$, we divide each individual measurement (e.g., $k_{i}+k_{j}$ for hierarchical attachement) by the sum over all available measurements (e.g., $p_{ij}=(k_{i}+k_{j})/\sum_{(i,j)\notin\mathcal{E}}(k_{i}+k_{j})$). Edges are then randomly selected based on these probability spaces $\{p_{ij}\}_{(i,j)\notin\mathcal{E}}$.
III-C Data
We obtained data for twenty large networks from two databases– the SNAP database [31] and the Network Repository [32]. While hierarchical complexity has been applied to several different types of networks (including social networks, protein networks and infrastructure networks) with mixed results compared to configuration models [15], it has yet to be applied to larger sized networks. Further, we have so far been unable to directly compare hierarchical complexity of networks of different sizes due to the lack of a normalisation.
Network were chosen to cover a wide range of sizes (1912-36692) and types (protein interaction networks, social networks, infrastructure networks, collaboration networks), and also to include groups of certain types of networks to explore relationships within and between network types. The number of nodes, edges and the network density for each network are shown in Table I.
Protein-protein interaction networks generated from co-expression correlations were taken from the Network Repository, which in turn derived these graphs from data from wormnet [33]. These were obtained for Homo Sapiens (HS), Derio Rario (DR)– zebrafish, Drosophilia Melanogaster (DM)– fruit fly, and caenorabditis elegans (CE)– a nematode. All of these are exceptionally well studied, model species for which the data is most extensive and reliable.
Also from the Network repository we took two infrastructure networks, one being the widely studied network of the Western States power grid of the US (power grid) [34] and the other being a network of international flights between airports where nodes are airports and edges are established where there are flights between those airports (open flights) [35]. All other networks were obtained from the SNAP repository.
We studied five collaboration networks within Physics disciplines, constructed from arXiv data. For these, edges are established between co-authors of papers. Topics are self-selected by authors during arXiv manuscript uploads. These topics are astrophysics (collab AstroPh), condensed matter physics (collab CondMat), general relativity (collab GrQc), high energy physics (collab HepPh), and high energy physics theory (collab HepTh).
We studied six social networks constructed from the twitch platform for six different languages– English (twitch ENGBE), French (twitch FR), German (twitch DE), Portuguese (twitch PTBR), Russian (twitch RU), and Spanish (twitch ES) [36]. Nodes are the users of twitch and edges are friendships between them. Facebook page-page is another social network where nodes represent official Facebook pages while edges are mutual likes between sites, collected through the Facebook Graph API in November 2017 and restricted to four categories of pages- politicians, governmental organisations, television shows and companies [36]. The LastFM social network is the network where nodes are users from Asian countries and edges are mutual follower relationships between them [37]. We also analysed a large email network between Enron employees (email Enron), originally made public by the Federal Energy Regulatory Commission during its investigation. Here, nodes are email addresses and edges are established wherever there are emails sent between addresses.
TABLE I: Statistics for twenty real world networks
Network $n$ $m$ $d$
email Enron 36692 367662 0.0005
Facebook 22470 171002 0.0007
collab CondMat 22167 186936 0.0008
collab AstroPh 16000 396160 0.0031
protein CE 15229 245952 0.0021
collab HepPh 12008 237010 0.0033
collab HepTh 9877 51971 0.0011
twitch DE 9498 153138 0.0034
lastFM Asia 7624 27806 0.0010
twitch ENGBE 7126 35324 0.0014
twitch FR 6549 112666 0.0053
collab GrQc 5242 28980 0.0021
power grid 4941 6594 0.0005
twitch ES 4648 59382 0.0055
protein HS 4413 108818 0.0112
twitch RU 4385 37304 0.0039
protein DM 4040 76717 0.0094
protein DR 3289 84940 0.0157
open flights 2939 30501 0.0071
twitch PTBR 1912 31299 0.0171
Allen brain model for use in large network experiment
For studying the effect of normalisation in very large graphs, we used the mouse V1 model from the Allen Institute for Brain Science [38]. This model contains approximately 230,000 neurons, and can be considered as a network where each neuron is a node and there is an edge between two nodes if they are connected by a synapse. We can sample geometric cylinders from this model using the associated code provided at [39], this allows us to construct networks that should be very similar structurally, but vary in size from anything up to 230,000 nodes. Due to computational limitations we only compute complexity values on up to 95,000 nodes.
Data for replication in attachment mechanism experiment
For studying the attachment mechanisms, we used a replication dataset from the ICON corpus consisting of 139 networks mostly describing biological, social and technological phenomena [40, 15]. These ranged from $n=50$ to $n=3155$ with a mean of 341 $\pm$ 462. Densities ranged from $d=0.0011$ to $d=0.3884$ with a mean of 0.0578 $\pm$ 0.0717.
IV Results
IV-A Normalisation results on random models
Fig. 2 shows the results of the normalisation applied to our chosen random graph models. For each random model (Erdös-Rényi, RGG, RHG, and RHGG) we generated 100 realisations with $n\sim U[50,10000]$ and $d\sim U[0,1]$ and plot results against both $n$ and $d$.
Notably, Erdös-Rényi random graphs have the lowest complexity of the models studied. Above this, the complexity of RGGs and RHGs is similar across all $n$ and $d$. The greatest complexity is clearly observed in the RHGGs, particularly at lower densities, indicating that greater complexity arises naturally through the combination of hierarchical and geometric structure.
While Erdös-Rényi random graphs, RGGs and RHGs have similar levels of complexity across density (highlighting the normalisation features of our measure across density), there is a different behaviour noted in RHGGs. Higher complexity at low densities compared to high densities in this instance can be considered a structural feature present in the graphs. We can understand very high densities as regarding the connectivity of the least important connections (specifically, the complement of the graphs) which in this instance are those between nodes with low degrees which are geometrically distant. It is reasonable to expect complexity here to be as low as for RHGs and RGGs.
As per the theoretical results for E-R random graphs, the experiments across the different models shows an inverse relationship for $\hat{R}$ with $n$ but little to no dependency on $d$. The reason for hierarchical complexity decreasing with increasing $n$ may be a true relationship of the complexity of these models as $n$ increases, rather than a normalisation issue. Indeed, we can expect that larger sample sizes of neighbourhood degree sequences given by larger random graphs would result in more homogeneous ordered sequences as they better approximate the global degree distributions.
Interestingly, for the RHGGs– which more accurately model real-world network structure– there is very little if any decrease with increasing $n$ for $n>1000$. This indicates that comparisons of $\hat{R}$ in large networks are reliable, but caution must still be taken when making comparisons in smaller networks.
Figure 2: Results show measurements for random realisations ($n\sim U[50,2500]$, $d\sim U[0,1]$) of different random graphs as denoted in the legend, against size $n$ and density $d$.
IV-B Normalisation results on increasing number of nodes
To extend our observations of the normalisation with respect to network size we studied the behaviour of the normalisation in very large graphs. In this case we use the mouse V1 model from the Allen Institute for Brain Science, see [38], and Erdös-Rényi random graphs, for which we know the normalised complexity value tends to 00 by Theorem 2.
We would expect networks of the same structure to have the same, or at least similar, HC. However, what do we mean by the “same structure”? If two networks have a different number of nodes they inherently have a different structure. In fact, the size of the network is related to the complexity, because as the size of random graphs increase their uniformity increases, this is due to the inverse relationship between variance and the sample mean. As such, we would expect that the normalised HC will decrease slightly as the number of nodes increases, but two sufficiently large networks of similar structure, but different size, would have similar complexity values.
Figure 3: The normalised hierarchical complexity of cylinders of increasing sizes of the Allen Brain V1 mouse model vs ER graphs of the same size and density.
The effect of this is well demonstrated in Fig 3. We see that the Erdös-Rényi random graph tends to 0 with increasing $n$, as expected. At the same time, for the Allen Brain for small $n$ the complexity is higher, but for larger $n$, roughly $n\geq 2000$, the complexity is very close between samples, appearing to tend to a non-zero limit in $n$. This demonstrates a behaviour expected of statistical complexity in dynamical systems, that randomness (aswell as regularity) vanishes to 0 complexity in the limit of $n$, while non-zero complexity is maintained in diverse structure [5].
IV-C Effect of degree heterogeneity on hierarchical complexity
Here we study the change in complexity as we increase heterogeneity among the RHGGs. We generated 1000 RHGGs of size $n=1000$ and $d\sim U[0,1]$. The heterogeneity was randomly determined in each iteration with $\sigma_{h}\sim U[0,1]$. Results are shown in Fig 4.
Figure 4: Top, average normalised hierarchical complexity per degree of 1000 heterogeneous geometric graphs with with $n=1000$, $d\sim U[0,1]$, $\sigma_{h}\sim U[0,1]$. Bottom, global normalised hierarchical complexity plotted against heterogeneity of these networks.
The results show that normalised HC has a consistent range of values across most degrees for heterogeneous geometric graphs, tapering off quickly at either end towards 1 and $n$. It is beneficial that the measure does not have a positive or negative relationship with degree so that the measure does not overly emphasise any particular range of degrees. Further, the fact the measure quickly tapers off towards very small degrees protects the measure from being influenced by the high levels of uncertainty of sampling at these small sample sizes.
HC is generally strongest in models with $\sigma_{h}$ between 0.2 and 0.4 and decreases towards low and high heterogeneity. When $\sigma_{h}$ is high the network becomes dominated by the hierarchical relationships, which should make the network more ordered. On the other hand, with low $\sigma_{h}$ the model gets closer to a random geometric graph which has low values of HC as shown in Fig 2. This is consistent with our expectations of HC being a statistical measure of complexity in networks.
IV-D Hierarchical complexity in large real-world networks
We compared unnormalised and normalised HC in the twenty large networks detailed in Table I. The results are shown in Table II. The values of the unnormalised HC range from 4 in the power grid network to 87858 in the German twitch network, a difference by a factor of nearly 22000. This range highlights the lack of utility of an unnormalised and unbounded measure for comparisons of different networks. On the other hand, the values of normalised HC range from 0.0022 in the power grid network up to 0.0830 in the Portugese twitch network, a difference by a factor of just under 38.
We can see that the normalised measure allows us to compare between these networks more clearly. For example, consider the twitch DE, twitch FR and twitch RU networks, which have the largest unnormalised HC values. These three networks are all constructed in the same way (friendships in twitch) and have similar densities (0.0034,0.0053, and 0.0039, respectively), and yet the twitch DE unnormalised HC value is 6 times bigger than the twitch RU unnormalised HC, suggesting that the twitch DE network is significantly more complex than the twitch RU network. However, when normalised these three networks all have very similar HC values, suggesting they have very similar levels of complexity, as one would expect.
Next consider the other three twitch networks: ES, PTBR and ENGBE. We can see that twitch ES and twitch ENGBE have very similar unnormalised HC values, but when normalised the twitch PTBR is twice as complex as twitch ES. This is not surprising considering they have markedly different densities, and we will see in the next section that even after normalisation the network density is correlated to the HC value in real world networks. We also see that the twitch ENGBE has much lower complexity than the other twitch networks, which is likely explained by the lower density of the network.
So we can see that our normalisation reduces the scale of the difference between complexities of networks, and allow us to better compare networks of different sizes. However, we also see that the density of the network still correlates with the normalised HC value.
TABLE II: Rankings of hierarchical complexity and normalised hierarchical complexity for twenty real world networks
$R$ Network $\hat{R}$ Network
87858 twitch DE 0.0830 twitch PTBR
35996 twitch FR 0.0684 twitch RU
14328 twitch RU 0.0612 twitch DE
11099 email Enron 0.0586 twitch FR
7625 twitch ES 0.0526 open flights
7167 twitch PTBR 0.0454 twitch ES
4216 twitch ENGBE 0.0360 protein DR
3275 facebook 0.0334 protein HS
2207 collab HepPh 0.0297 email Enron
1283 protein HS 0.0279 protein DM
1015 protein DR 0.0272 twitch ENGBE
873 collab AstroPh 0.0258 collab HepPh
755 protein CE 0.0182 facebook
752 protein DM 0.0154 collab GrQc
582 open flights 0.0147 lastFM Asia
293 LastFM Asia 0.0124 protein CE
261 collab CondMat 0.0105 collab AstroPh
68 collab GrQc 0.0056 collab HepTh
39 collab HepTh 0.0054 collab CondMat
4 power grid 0.0022 power grid
IV-E Growth of hierarchical complexity in real world networks
We found that hierarchical complexity was positively correlated with network density in the twenty networks from Table I ($\rho=0.7203$, $p=0.0005$), Fig. 5, top-left. We confirmed this observation with a network dataset of 139 smaller networks ($n\in[50,3155]$, $\rho=0.5325$, $p=1.5\times 10^{-11}$) [40, 15], Fig. 5, top-right. To discount the potential confounding effect of network size, $n$, on these correlations, we implemented linear regression on network density with network size as a predictor and found that the correlations of the residuals of the regression with HC were still significant in both cases– $\rho=0.6526$, $p=0.0023$ for the 20 large networks and $\rho=0.2294$, $p=0.0066$ for the 139 small-to-medium sized networks.
Figure 5: Top row shows scatterplots visualising the positive association between density and hierarchical complexity in real-world networks. Spearman’s correlation coefficient and associated $p$-value shown inset. Bottom row shows average results of the values of HC as we increase density of networks according to the attachment mechanisms as described in the legend.
At the same time, we have shown that hierarchical complexity shows strong normalisation with respect to density for many types of graph. The relationship between density and hierarchical complexity in real world networks is therefore unlikely due to a lack of normalisation, but is a true relationship requiring a mechanistic explanation. To try to explain this relationship we applied several attachment algorithms, as described in section III.B, to real-world networks to artificially increase their density and see if any would consistently lead to the targeted increase in HC.
Starting from the original networks, we increased density according to the four different attachment mechanisms in Section III.B from 0.1% up to a 2% increase in density in steps of 0.1%. Normalised HC was recorded at each step. The average results across the network datasets are shown in Fig. 5, bottom row. Patterns replicate in both datasets. Most evidently, none of the attachment mechanisms could explain an increase in complexity with increasing density as seen across the real-world networks. If anything, all mechanisms showed a decrease in complexity as density was increased.
Notably, the hierarchical & similar attachment mechanism which promotes connections primarily among those with most connections and those with greatest overlap of connections, did exhibit the least decrease in both datasets, as may be expected from our results in the RHGG models, but it still was not able to explain an increase in complexity.
V Discussion
Our modelling demonstrated greater statistical complexity arising through the combination of hierarchical and geometric components. While the space of random geometric graphs is regular, the nodes being placed randomly in that space opens up pockets of higher and lower connectivity, as quantified through degree correlations [27]. This may be understood as a classic instance where randomness and regularity interact to generate some degree of complexity. Similarly, heterogeneous random graphs contain an ordered structure in terms of the hierarchy of node degrees, combined with randomness of connections established through the configuration model procedure. By giving a randomly allocated log-normal node fitness to the nodes randomly placed in Euclidean space we see an amplification in terms of complexity. This is particularly interesting since heterogeneous geometric networks closely model many aspects of real-world networks [20]. Future work will explore whether access to heterogeneous connectivity patterns in networks facilitates the formation of more heterogeneous functionality and therefore assess the advantages conferred by the combination of hierarchy and geometry which real-world networks appear to incorporate almost universally.
In the results of HC among the 20 large real-world networks some patterns appear. The general trends among the three groups of networks we have (6 twitch networks, 4 protein networks, 5 physics collaboration networks) indicate that the twitch networks tend to have highest complexity, while protein networks have fairly high complexity and physics collaboration networks have low complexity. The power grid having the lowest complexity can be expected as it is a network with high geometrical constraints and we might expect some specific universal design principles in its construction. On the other hand, online social networks of twitch are largely free from geometrical constraints and may reflect the diversity of social relationships. However, sample sizes would need to be increased to provide stronger evidence for any such generalisations.
The results from the attachment mechanism experiments showed that explanations for the positive relationship between density and hierarchical complexity were not the result of preferential attachment mechanisms for hierarchy or similarity of nodes, nor the combination of the two. This shows us that the increase in HC with respect to density is not a trivial observation. Going further, it points towards a preference of real-world networks towards complexity in novel attachments.
The evidence that random geometric graphs have non-zero HC in the thermodynamic limit gives an interesting insight into the geometrical nature of HC. It tentatively points towards a definitive notion of statistical complexity of networks. Essentially, if we take seriously the established notion of networks as embedded in a latent geometrical space [41, 19], then we can start to conceptualise measures of statistical complexity of networks such as HC as attempting to measure the irregularity of the distribution of points over that space.
In the future, we intend to thoroughly explore the application of normalised HC for analysis of brain networks, across scales, across different types of networks, and eventually, across species. We will also explore applications to protein-protein interaction networks utilising the vast dataset available through SNAP [31] in order to begin to answer questions regarding the relationship between HC of protein-protein interaction networks and evolutionary parameters. Applications to social networks hold obvious appeal given the consistently high complexity we noticed in the twitch social networks. Generally speaking, our understanding of statistical complexity of networks in different fields is limited by a lack of parsimonious measurements for its quantification. The tools offered here can therefore help researchers to begin to answer fundamental questions regarding the existence and extent of statistical complexity of real-world networks, especially in light of the larger and higher quality datasets becoming evermore available.
VI Conclusion
We proposed and demonstrated the utility of a normalisation for hierarchical complexity. We proved that this measure is bounded above by $2$ and tends to zero for Erdös-Rényi random graphs with increasing size. This is analogous to a defining characteristic required of a statistical complexity measure in dynamical systems. We then demonstrated that, while random graph models containing hierarchy and geometry individually had lower complexity, the combined components of hierarchy and geometry is enough to create HC of a similar level to real-world networks. However, real-world networks displayed an association between HC and density that could not be explained solely by these components. Likewise, modelling attachment mechanisms following hierarchy, geometry, or their combination could not explain this association. We conjecture that real-world networks therefore have in-part a preference for attachments which increase complexity of the interacting system. We provide a parsimonious measure for statistical complexity of networks which is ready to be applied to answering questions and gathering new insights into the degrees of complexity in various fields such as neuroscience, protein biology and social networks.
Acknowledgment
The authors would like to thank Dr April S. Kleppe for her thoughtful feedback on our manuscript.